Workers: Akutsu, K. Nagano, Aritomi K. Tanaka, Kita
- Delivered an oplev pylon.
- Delivered a K400 flange (VN7877 by Mirapro), which had been washed at Mirapro.
Checked the polarity of LVDTs and actuator coils, and compared with the one measured in Mitaka shown here.
When we took up the side of the table where LVDT-V1 was installed, the raw LVDT-V1 signal (VIS-TMSX_LVDTINF_**_INMON) showed negative signal; in the same manner, V2-> positive, V3-> negative, and all the horizontal signals (H1,H2,H3) were positive; they were consistent with those of in Mitaka.
When we input plus (about +3000 counts) signal to actuator coil H1, the raw LVDT H1, H2, and H3 (VIS-TMSX_LVDTINF_**_INMON) showed all positive; in the same manner, act coil H2 positive -> all LVDT-Hx positive, act coil H3 positive -> all LVDT-Hx positive, act coil V1 positive (+5000 counts) -> LVDT-V1 negative, act coil V2 positive -> LVDT-V2 positive, act coil V3 positive -> LVDT-V3 negative.
To translate the results of the actuator coils, considering the polarity of the raw LVDT signals, all the actuator coils were set to actuate the table to the positive directions when we input positive counts to them; the positive directions here followed the difinitions in here; positive yaw is counter clock wise from the top view, and upward is positive vertical. Note that we have replaced all the horizontal actuator coils in Kamioka , so it would be natural if their polarity are not consistent with those in Mitaka.
- Checked roughly the direction of the collimated green ouput from the 2nd BRT lens; it seemed the beam (1) rose up during it travels to the Xend, (2) tilted in +Y arm direction in yaw. We will check them in detail later.
- Align and balance the BRT and the TMS-VIS.
- Configure the digital models properly.
Clean the sides of the optics to be installed into the vacuum chamber. (Ethanol+Acetone wiping with vectra alpha, then First Contact.) See here.
- Detach the remaining white tapes just before closing the chamber; be noted that there are white tapes on the PEEK connectors at the inside of the flanges.
- Torque completely the bolts which are fixing the upper and the lower base plates.
- Insert metal gaskets before torquing the ICF203 flanges with Dsub connectors.
- Check horizontal actuation
- Calibrate LVDT again
- Put calibration factor of Geophones
- Install Tome-Kanagu for Dsub in DAC0 port #4
- Put tags to cables
- Align BRT
- Start to install GPT part
- Damping of TMS-VIS
- Tester!! lend from DGS or PEM??
- M8 eyebolts for managing the dummy weight plates...
- First contact to clean the side diameter of the optics.
- Need a signal generator for LVDTs
- A document for learning how to release or fix the suspended tables (by Uraguchi?).
- Cable tags
- IR mirrors (4 PYD-20, 3 PYD-10)
- Tome-Kanagu for Dsub
- Power taps
- ONE DAC PORT!!!!